y=x-5
What is the x and y intercepts
: y=x-5
What is the x and y intercepts

Answer by Alan3354(1926)

y=x-5
What is the x and y intercepts?
------------
To find the x-intercept, set y = 0
To find the y-intercept, set x = 0
-------------------
y = x+5
@ x = 0, y =5
The point is (0,5) y-intercept
--------------
y = x+5
@ y = 0, x = -5
The point is (-5,0) x-intercept